OpenStudy (christos):

How do we determine if a function is increasing or decresing just by testing its second derivative ?

OpenStudy (anonymous):

Sorry I dont know I am only in 7th grade lol sorry :(

OpenStudy (christos):

its np

OpenStudy (misty1212):

HI!!!

OpenStudy (misty1212):

you want the first and second derivative right?

OpenStudy (misty1212):

just checking it?

OpenStudy (christos):

well I know how to test the fist derivate but how about the second ?

OpenStudy (christos):

can I get info by just testing the second derivative ?

OpenStudy (misty1212):

the second derivative indicates concavity you need that?

OpenStudy (christos):

I wanna test increasing decreasing

OpenStudy (christos):

if there was any chance I could derive that from concativity ?

OpenStudy (misty1212):

it does not tell you if the function is increasing or decreasing the first derivative does that, and your answer is correct

OpenStudy (christos):

without using first derivative tests

OpenStudy (misty1212):

you cannot

OpenStudy (misty1212):

you can check if a critical point is a local max or min using the second derivative, but you cannot use the second derivative to see if the function is increasing or decreasing i can give you an easy example if you like

OpenStudy (misty1212):

think of \(f(x)=x^2\) which is decreasing on \((-\infty,0)\) and increasing on \((0,\infty)\) the second derivative is \(f''(x)=2\)

OpenStudy (christos):

can I test if a function has a rel max/min with just first derivative tests ?

OpenStudy (misty1212):

yes, you can you can use the second derivative, or the first derivative the check is different

OpenStudy (christos):

can you also please tell me whats the intuitive difference between a stationary point and a critical point ?

OpenStudy (misty1212):

we can use your example and check \(\sqrt2\) if you like to see if it is a max or min using the first derivative, and check again using the second derivative

OpenStudy (christos):

yes pleae

OpenStudy (christos):

"we can use your example and check 2√ if you like to see if it is a max or min using the first derivative, and check again using the second derivative " Let's do that

OpenStudy (misty1212):

ok

OpenStudy (misty1212):

you have decided correctly that \(f\) is increasing on \((-\sqrt2,\sqrt2)\) and then decreasing on \((\sqrt2,\infty)\)

OpenStudy (misty1212):

in other words, the derivative is positive on the first interval and then becomes negative on the second interval that means that since it changes sign, from positive to negative at \(\sqrt2\), \(\sqrt2\) must give a local max

OpenStudy (misty1212):

a mental picture shows it going up, then going down 

OpenStudy (christos):

is that a stationary , critical point or both ? ?

OpenStudy (misty1212):

ok now lets answer that question

OpenStudy (misty1212):

the distinction is a stationary point \(a\) is a number with \(f'(a)=0\)

OpenStudy (misty1212):

whereas a critical point is either \(f'(a)=0\) or \(f'(a)\) does not exist

OpenStudy (christos):

so rel max/min can only be criticals

OpenStudy (misty1212):

so all stationary points are critical points, but no vice versa would you like a simple example?

OpenStudy (christos):

is my assumption that "all rel max/min can only be criticals and not stationary" correct ?

OpenStudy (misty1212):

no

OpenStudy (misty1212):

ALL stationary points are also critical points an example may help

OpenStudy (christos):

so then what's the intuition behind the dinstinction of criticals and stationary

OpenStudy (christos):

on a graph

OpenStudy (christos):

graphicly *

OpenStudy (misty1212):

ok i will draw two pictures both of which have 0 as a critical point, one being stationary, the other not stationary

OpenStudy (misty1212):

the functions i will draw are very simple one is \(x^2\) the other \(|x|\)

OpenStudy (misty1212):